# arXiv:2602.22437v2[cs.DC] 27 Feb 2026

### v2

@@ 5 Grouped RaggedShard for Performance
The core idea is a case analysis of how each tensor aligns with shard boundaries in any valid layout: (1) it lies entirely within a single local shard; (2) it straddles two adjacent shards, but doesn’t contain a full shard; (3) it fully contains at least one shard. If every tensor falls into cases (1)–(2), feasibility is monotone in the shard size $S$: whenever a layout exists for $S$, it also exists for $S+\Delta$ (where $\Delta$ is the base alignment quantum). Because every shard includes an inter-tensor boundary, we can always absorb additional $\Delta$ as padding. If any tensor is in case (3), the feasible shard sizes must be multiples of $L=\operatorname{LCM}\{\,g_{t}\mid\text{$t$ is in case (3)}\,\}$. In this regime, feasibility is monotone over multiples: if $kL$ is feasible, then (k+1)L(k{+}1)L is also feasible. We therefore enumerate the case-(3) set and binary-search the minimal $S$ over the corresponding multiples, as shown in Line 17-21. To avoid exponentially enumerating all Case-(3) subsets, we sort tensors by element count and enumerate only the granularity prefixes, yielding a 2-approximation. Inside the feasibility checker, we define $dp(t,i)$ as the minimum number of devices (shards) required to store all atomic units up to the $i$-th unit of tensor $t$. Although the index space is as large as numerical space, $dp(t,i)$ is monotone within a tensor: $dp(t,i)\leq dp(t,i{+}1)$. So within each tensor there are at most $m$ distinct values. In Line 8-9, we exploit this by batching contiguous indices into segments and skipping dp(t,)dp(t,*) evaluations at intermediate indices, achieving the stated time complexity.
@@
Figure 7: Distributed Buffer (DBuffer) for high performance communication. A 2D DBuffer for AllGather’ing parameters is shown; Reversely, a 2D DBuffer redistributing from (Partial, Partial) to (Replicate, Shard) implements 2D gradient reduction with ReduceScatter and AllReduce.
@@
Distributed Buffer (DBuffer)   Beyond grouping RaggedShard DTensors, the underlying communication buffer also plays a vital role in achieving high performance of communication, computation, and memory efficiency. To this end, veScale-FSDP proposes a new primitive, Distributed Buffer (DBuffer), to optimize the performance of grouped DTensors. Figure 7 shows the design. First, inspired by DTensor, DBuffer provides global buffer semantics over an $N$-dimensional device topology, with a sharding specification along each dimension, abstracting away the complexity of $N$-D communication and operations. Second, DBuffer takes a group of tensors and executes group-level operators rather than per-tensor operators. For example, before communication, each tensor may need to launch its own CUDA kernels for add, scale, zero, or copy (which may differ across tensors), incurring fragmented compute overhead and blocking communication. With DBuffer, identical kernels across tensors are fused before communication, reducing blocking time. Third, DBuffer offers zero-copy before and after communication by leveraging RaggedShard’s planning algorithm and providing a persistent address mapping to each tensor’s data pointer, minimizing memory footprint and fragmentation. Lastly, DBuffer uses in-place communication and computation.
@@ 6 Evaluation
  • How does veScale-FSDP planner minimize padding, and what is the algorithm overhead (§6.4)?
@@
[…]
@@ 6.1 End-to-End Performance
Figure 8 compares the performance of veScale-FSDP to baselines on three representative models on 1024 GPUs.
@@
Throughput: on the MoE models, veScale-FSDP is $1\sim$66% faster than all baselines. On LLaMA-3-70B veScale-FSDP is 5% faster than DeepSpeed, FSDP1, and FSDP2, and slightly ahead of Megatron-FSDP. The higher throughput arises from optimized communication overlapping, DBuffer-based zero-copy collectives, and flexible sharding granularities that avoid padding overhead. In contrast, DeepSpeed emits fragmented collectives [deepspeed_AG], while FSDP1 exhibits communication bubbles where data movement operations block NCCL progress, under-utilizing the network in both systems. FSDP2 relies on the per-parameter DTensor even-sharding format that introduces interleaved copy-out after all-gather and interleaved copy-in before reduce-scatter; together these copies can consume up to 14% of a training iteration and hence reduce throughput. In addition, FSDP1 and FSDP2 overlook NCCL address alignment caveat, leading to substantial degenerate communication performance in certain cases [wu2025terabyte]. Although Megatron optimized for zero-copy collectives, its fixed Stride(0) sharding granularity, where to remain consistent with upstream DTensor Shard(0) semantic for distributed checkpointing, induces 33% buffer padding inflation in MOE models thus slows the collectives [megatron_fsdp]. Our experiments show that veScale-FSDP achieves linear scalability; detailed analysis appears in §6.2.
@@
Memory: Across benchmarks, veScale-FSDP reduces peak reserved memory by 16–30%. The memory saving stems from deterministic, batched DBuffer memory management: we explicitly manage stream dependencies for predictable memory deallocation, and we batch allocations to reduce fragmentation. By contrast, DeepSpeed and FSDP1 inherit non-deterministic deallocations from implicit record_stream [per_parameter_shard_rfc], which often prevents the caching allocator from reusing buffers, inflating peak reserved memory by 20%. Relative to FSDP2 ’s per-parameter eager allocation, our batched policy yields a further 12% reduction. Megatron’s padding-inflated buffers not only degrade collective efficiency but also raise peak memory by 33% in MoE experiments; its mixed-precision support persists low-precision buffers, consuming 24% more memory than veScale-FSDP in the LLaMA-3 experiments. Lower reserved memory translates directly into higher end-to-end efficiency: under high memory pressure, the PyTorch caching allocator issues device frees that synchronize with the driver and stall training. In terms of scalability, veScale-FSDP ’s memory footprint decreases monotonically as the FSDP group size increases and grows only marginally with the replication factor, matching scaling expectations. A notable exception appears with GPT-OSS: FSDP2 trains at 128 devices but OOMs at 256. The per-parameter sharding design in DTensor requires padding to enforce even splits along the sharded dimension; with 128 experts spread over 256 devices, the all-gather buffer effectively doubles, exhausting memory. While FSDP2 allows custom sharding along other dimensions, it requires manual padding and thus doubles the interleaved-copy overhead, making it prohibitively expensive (recall Table 1).
@@ 6.2 Scalability and Composability
Figure 9: Scalability of veScale-FSDP in up to 10K GPUs and 2.4T model size.
The flexibility of RaggedShard also enables seamless integration with complementary parallelization strategies such as expert parallelism (EP) [lepikhin2020gshard]. Combining these techniques allows veScale-FSDP to efficiently scale training to internal models with up to 2.4T parameters on as many as 10K GPUs, as shown in Figure 9. Note that we evaluate scalability of MoE, because MoE workloads are often more challenging to scale under FSDP: sparse expert computation lowers per-GPU compute while requires substantial all-gather/reduce-scatter traffic, making communication and padding overheads more significant.
@@
We show the flexibility of RaggedShard DTensor using two examples: 8-bit Adam and distributed Muon optimizer.
8-bit Adam optimizer. 8-bit Adam [dettmers8] applies block-wise INT8 quantization to the gradient statistics, substantially reducing optimizer-state memory. To enable 8-bit Adam, veScale-FSDP exposes orig_param_policy interface that lets users set the quantization granularity per parameter. In our setup, we use $32\times 32$ blocks and assign matrix parameters to 32-row block granularity. With this layout, each device quantizes its local shard independently without any communication, and block boundaries are perfectly preserved by RaggedShard. In contrast, existing FSDP systems do not natively track such block boundaries, so enabling block-wise 8-bit Adam often requires intrusive system changes or manual collectives to exchange quantization metadata, incurring both complexity and overhead.
@@
We implement the 8-bit Adam using veScale-FSDP with few lines of code and provide the evaluation in Figure LABEL:fig:adam_loss. We compare 8-bit Adam under distributed data parallelism (DDP) and veScale-FSDP. The loss curves track closely, with occasional spikes characteristic of reduced-precision optimizer states. The small difference stems from the gradient-reduction schedule: DDP uses bucketed all-reduce, whereas veScale-FSDP performs layer-wise reduce-scatter. (Note that the loss curves in Figure 10 are not directly comparable: for 8-bit Adam we use a smaller learning rate to mitigate overflow/underflow in reduced precision.)
@@
4:  // Choose compute device via load balancing
5:  $r\leftarrow\textsc{SelectRoot}()$
6:  // Unshard to root via redistribute
7:  $\mathbf{o}\leftarrow\textsc{Redistribute}(\mathbf{u},{\small\texttt{RaggedShard}}(r))$
8:  // Muon update: Newton–Schulz on full tensor.
9:  $\mathbf{o}\leftarrow\textsc{NewtonSchulz}(\mathbf{o})$
10:  // Redistribute update back.
11:  $\mathbf{o}\leftarrow\textsc{Redistribute}(\mathbf{o},\mathrm{placement}(\mathbf{u}))$
12:  $\mathbf{w}\leftarrow\mathbf{w}-\eta\,\mathbf{o}$
13:end for
Distributed Muon optimizer. The matrix-sign preconditioner (e.g., Newton–Schulz) of Muon requires the full 2D parameter matrix with its original shape. Algorithm 2 sketches the distributed Muon optimizer enabled by RaggedShard. Thanks to RaggedShard’s capability to support uneven sharding, users can write Muon’s parameter-gather step in a clean SPMD way: after redistribution, only the root rank holds the full 2D parameter, so the Newton–Schulz update becomes a no-op on other ranks. As lines 4–7 show, the algorithm selects a root via load balancing and unshards to it using the standard DTensor redistribute with RaggedShard placement. Lines 8–9 run the Muon matrix iteration only on the root that holds the full tensor. Finally, lines 10–12 redistribute the update back to the original device and apply it. Therefore, users do not need to handle the complex logic of communication and can further overlap communication with computation via asynchronous redistribute. In addition, our optimized Muon reaches 47.3% MFU on 256 GPUs by exploiting the communication-computation overlapping and using torch.compile to further increase compute density.
@@ 6.4 Planning Quality
A major design objective of the planning algorithm (Algorithm 1) is to enable arbitrary granularity of RaggedShard, while minimizing padding overhead and thus reducing communication volume. The quality of planning algorithm can be directly evaluated by padding size. We evaluate it by benchmarking DeepSeek-v3-671B and GPT-OSS-120B across varying device counts. Following the DeepSeek-style quantization scheme, we quantize only the FFN weights (most parameters) and sweep the row granularity of expert-MLP matrices over 128, 16, 1. The 128-row setting reproduces DeepSeek’s $\times128$ tiling (i.e., weights can be sliced into $\times128$ blocks). We then report the resulting relative padding ratios and analyze the root cause of the extra padding.
@@
Lastly, we also evaluate the overhead of planning algorithm itself: the algorithm runtime is less than 0.3 seconds across all experiments, which is one-time and negligible in distributed training initialization.
@@ 6.5 Performance Breakdown
Table 2: Component ablation for 8-bit Adam, in terms of normalized throughput when disabling each component. N/A means impossible without intrusively changing model/optimizer code or managing custom collectives.
@@
  • DBuffer. Disabling DBuffer drops throughput by 7.2%, reflecting the copy-in/copy-out overhead around collectives when communication buffers require copy.
@@
  • RaggedShard DTensor. Disabling RaggedShard makes it effectively non-runnable: users must either (i) carefully change every model and optimizer tensor so that $32\times 32$ block boundaries align with shard boundaries, or (ii) manually implement complex collectives (e.g., per-block metadata exchange and state gathering) to recover block-wise semantics. We therefore report this setting as N/A to indicate it’s not meaningfully usable.
@@ 7 Lessons Learned
During the deployment of veScale-FSDP for real industrial workloads that use more than 10K GPUs, we summarize the key lessons we have learned in the following.
@@
Lesson-2: Design system abstractions on the shoulders of giants. DTensor provides a powerful abstraction that already supports a wide range of parallelization techniques. By designing new abstractions on top of DTensor, we can seamlessly integrate existing parallelization strategies. In our work, RaggedShard placement is implemented as an optional placement on DTensor, enabling easy collaboration of established infrastructure such as tensor and expert parallelism, as well as mature training tools like distributed checkpointing [dcp]. This approach minimizes engineering effort while contributing to a shared ecosystem that benefits the broader community. In fact, RaggedShard has already appeared as a planned feature on the official roadmap [pytorch_2026_h1_roadmap] of PyTorch.

### v3

@@ 5 Grouped RaggedShard for Performance
The core idea is a case analysis of how each tensor aligns with shard boundaries in any valid layout: (1) it lies entirely within a single local shard; (2) it straddles two adjacent shards but does not fully contain a shard; and (3) it fully contains at least one shard. If every tensor falls into cases (1) or (2), feasibility is monotonic in the shard size $S$: whenever a layout exists for $S$, it also exists for $S+\Delta$, where $\Delta$ is the smallest alignment unit required by the layout. Because every shard includes an inter-tensor boundary, the additional $\Delta$ can always be absorbed as padding. If any tensor falls into case (3), the feasible shard sizes must be multiples of $L=\operatorname{LCM}\{\,g_{t}\mid\text{$t$ is in case (3)}\,\}$. In this regime, feasibility is monotonic over multiples of $L$: if $kL$ is feasible, then $(k+1)L$ is also feasible. We therefore binary-search for the minimal feasible $S$ over the corresponding multiples, as shown in Lines 21–25. To avoid enumerating an exponential number of candidate case-(3) sets, we sort tensors by element count and consider only prefixes of this sorted order, yielding a 2-approximation. Within CheckValidShard, we define $dp(t,i)$ as the minimum number of devices (shards) required to store all previous tensors and the first $i$ sharding blocks of tensor $t$. Although the DP state space indexes every block position within a tensor, $dp(t,i)$ is monotonic: $dp(t,i)\leq dp(t,i+1)$. Therefore, each tensor has at most $m$ distinct DP values. In Lines 10–13, we exploit this property by grouping contiguous indices into segments and skipping intermediate $dp(t,\ast)$ calculations, which yields the stated time complexity.
@@
Figure 7: Distributed Buffer (DBuffer) for high-performance communication. A 2D DBuffer for parameter AllGather is shown; conversely, a 2D DBuffer redistributing from (Partial, Partial) to (Replicate, Shard) implements 2D gradient reduction with ReduceScatter and AllReduce.
@@
Distributed Buffer (DBuffer)   Beyond grouping RaggedShard DTensors, the underlying communication buffer also plays a vital role in achieving high communication, computation, and memory efficiency. veScale-FSDP introduces a new primitive, Distributed Buffer (DBuffer), for efficient grouped DTensors. Figure 7 shows the design. First, inspired by DTensor, DBuffer provides global buffer semantics over an $N$-dimensional device topology, with a sharding specification along each dimension, abstracting away the complexity of $N$-D communication and operations. Second, DBuffer takes a group of tensors and executes group-level operators rather than per-tensor operators. For example, before communication, each tensor may need to launch its own CUDA kernels for add, scale, zero, or copy (which may differ across tensors), incurring fragmented compute overhead and blocking communication. With DBuffer, identical kernels across tensors are fused before communication, reducing blocking time. Third, DBuffer offers zero-copy access before and after communication by leveraging RaggedShard’s planning algorithm and providing a persistent address mapping to each tensor’s data pointer, thereby minimizing memory footprint and fragmentation. Finally, DBuffer uses in-place communication and computation.
@@ 6 Evaluation
  • How does the veScale-FSDP planner minimize padding, and what is the algorithm overhead (§6.4)?
@@
Hardware: We ran the experiments §6.1, §6.4, and §6.5 on a NVIDIA H800 cluster; each node contains 8×H800 GPUs (979 BF16 TFLOPS, 80 GB HBM) and 400 GB/s NVLinks. The experiments §6.2 and §6.3 were conducted on a NVIDIA Hopper cluster.
@@ 6.1 End-to-End Performance
Figure 8 compares the performance of veScale-FSDP against the baselines on the three representative models introduced in the previous section with 1024 GPUs.
@@
Throughput: On the MoE models, veScale-FSDP is $1\sim$66% faster than all baselines. On LLaMA-3-70B, veScale-FSDP is 5% faster than DeepSpeed, FSDP1, and FSDP2, and slightly ahead of Megatron-FSDP. The higher throughput arises from optimized communication overlapping, DBuffer-based zero-copy collectives, and flexible sharding granularities that avoid padding overhead. In contrast, DeepSpeed emits fragmented collectives [deepspeed_AG], while FSDP1 exhibits communication bubbles where data movement operations block NCCL progress, under-utilizing the network in both systems. FSDP2 relies on the per-parameter DTensor even-sharding format that introduces interleaved Copy-Out after AllGather and interleaved Copy-In before ReduceScatter; together these copies can consume up to 14% of a training iteration and hence reduce throughput. In addition, FSDP1 and FSDP2 do not enforce NCCL buffer address alignment, so collectives may operate on unaligned memory addresses and suffer substantial communication performance degradation in certain cases [wu2025terabyte]. Although Megatron is optimized for zero-copy collectives, its fixed Shard(0) sharding granularity, chosen to remain consistent with the upstream DTensor Shard(0) semantics for distributed checkpointing, induces 33% buffer padding inflation in MoE models and thus slows collective communication [megatron_fsdp]. Our experiments show that veScale-FSDP achieves linear scalability; detailed analysis appears in §6.2.
@@
Memory: Across benchmarks, veScale-FSDP reduces peak reserved memory by 16–30%. The memory saving stems from deterministic, batched DBuffer memory management: We explicitly manage stream dependencies for predictable memory deallocation, and we batch allocations to reduce fragmentation. By contrast, DeepSpeed and FSDP1 inherit non-deterministic deallocations from PyTorch’s implicit record_stream mechanism [per_parameter_shard_rfc], which often prevents the caching allocator from reusing buffers, inflating peak reserved memory by 20%. Relative to FSDP2 ’s per-parameter eager allocation, our batched policy yields a further 12% reduction. Megatron’s padding-inflated buffers not only degrade collective efficiency but also raise peak memory by 33% in MoE experiments; its mixed-precision support persists low-precision buffers, consuming 24% more memory than veScale-FSDP in the LLaMA-3 experiments. Lower reserved memory translates directly into higher end-to-end efficiency: under high memory pressure, the PyTorch caching allocator issues device frees that synchronize with the driver and stall training. In terms of scalability, veScale-FSDP’s memory footprint decreases monotonically as the FSDP group size increases and grows only marginally with the replication factor, matching scaling expectations. A notable exception appears with GPT-OSS: FSDP2 trains at 128 devices but OOMs at 256. The per-parameter sharding design in DTensor requires padding to enforce even splits along the sharded dimension; with 128 experts spread over 256 devices, the AllGather buffer effectively doubles, exhausting memory. While FSDP2 allows custom sharding along other dimensions, it requires manual padding and thus doubles the interleaved-copy overhead, making it prohibitively expensive (recall Table 1).
@@ 6.2 Scalability and Composability
Figure 9: Scalability of veScale-FSDP. (a) Weak scaling to 8K GPUs. (b) Strong scaling to 8K GPUs. (c) Strong scaling at larger scale with normalized throughput and normalized GPU count. (d) Model scaling on 1K GPUs up to 2.4T parameters.
The flexibility of RaggedShard also enables seamless integration with complementary parallelization strategies such as Expert Parallelism (EP) [lepikhin2020gshard]. Combining these techniques allows veScale-FSDP to efficiently scale training to internal models with up to 2.4T parameters on as many as 10K Hopper GPUs, as shown in Figure 9. Note that we evaluate scalability of MoE, because MoE workloads are often more challenging to scale under FSDP: sparse expert computation lowers per-GPU compute while requiring substantial AllGather/ReduceScatter traffic, making communication and padding overheads more significant.
@@
We demonstrate the flexibility of RaggedShard DTensor using two examples: the 8-bit Adam optimizer and the distributed Muon optimizer.
8-bit Adam optimizer. 8-bit Adam [dettmers8] applies block-wise INT8 quantization to the gradient statistics, substantially reducing optimizer-state memory. To enable 8-bit Adam, veScale-FSDP exposes an orig_param_policy interface that lets users set the quantization granularity per parameter. In our setup, we use $32\times 32$ blocks and assign matrix parameters to 32-row block granularity. With this layout, each device quantizes its local shard independently without any communication, and block boundaries are perfectly preserved by RaggedShard. In contrast, existing FSDP systems do not natively track such block boundaries, so enabling block-wise 8-bit Adam often requires intrusive system changes or manual collectives to exchange quantization metadata, incurring both complexity and overhead.
@@
We implement 8-bit Adam using veScale-FSDP with few lines of code and provide the evaluation in Figure LABEL:fig:adam_loss. We compare 8-bit Adam under distributed data parallelism (DDP) and with veScale-FSDP. The loss curves track closely, with occasional spikes characteristic of reduced-precision optimizer states. The small difference stems from the gradient-reduction schedule: DDP uses bucketed AllReduce, whereas veScale-FSDP performs layer-wise ReduceScatter. (Note that the loss curves in Figure 10 are not directly comparable: for 8-bit Adam, we use a smaller learning rate to mitigate overflow/underflow in reduced precision.)
@@
4:  $p\leftarrow\mathrm{placement}(\mathbf{u})$ // original DTensor placement
5:  // Choose compute device via load balancing
6:  $r\leftarrow\textsc{SelectRoot}()$
7:  // Unshard to root via redistribute
8:  $\mathbf{o}\leftarrow\textsc{Redistribute}(\mathbf{u},{\small\texttt{RaggedShard}}(r))$
9:  // Muon update: Newton–Schulz on full tensor.
10:  $\mathbf{o}\leftarrow\textsc{NewtonSchulz}(\mathbf{o})$
11:  // Redistribute update back.
12:  $\mathbf{o}\leftarrow\textsc{Redistribute}(\mathbf{o},p)$
13:  $\mathbf{w}\leftarrow\mathbf{w}-\eta\,\mathbf{o}$
14:end for
Distributed Muon optimizer. The matrix-sign preconditioner (e.g., Newton–Schulz) of Muon requires the full 2D parameter matrix with its original shape. Algorithm 2 sketches the distributed Muon optimizer enabled by RaggedShard. Thanks to RaggedShard’s capability to support uneven sharding, users can write Muon’s parameter-gather step in a clean SPMD way: after redistribution, only the root rank holds the full 2D parameter, so the Newton–Schulz update becomes a no-op on other ranks. As lines 5–8 show, the algorithm selects a root via load balancing and unshards to it using the standard DTensor redistribute with RaggedShard placement. Lines 9–10 run the Muon matrix iteration only on the root that holds the full tensor. Finally, lines 11–13 redistribute the update back to the original device and apply it. Therefore, users do not need to handle the complex logic of communication and can further overlap communication with computation via asynchronous redistribute. In addition, our optimized Muon reaches 47.3% MFU on 256 Hopper GPUs by exploiting the communication-computation overlapping and using torch.compile to further increase compute density.
@@ 6.4 Planning Quality
A major design objective of the planning algorithm (Algorithm 1) is to enable arbitrary granularity of RaggedShard, while minimizing padding overhead and thus reducing communication volume. The quality of the planning algorithm can be directly evaluated by padding size. We evaluate it by benchmarking DeepSeek-V3-671B and GPT-OSS-120B across varying device counts. Following a DeepSeek-style quantization scheme, we quantize only the FFN weights (most parameters) and sweep the row granularity of expert-MLP matrices over 128, 16, 1. The 128-row setting reproduces DeepSeek’s $\times128$ tiling (i.e., weights can be sliced into $\times128$ blocks). We then report the resulting relative padding ratios and analyze the root cause of the extra padding.
@@
We also evaluate the overhead of the planning algorithm itself: the algorithm runtime is less than 0.3 seconds across all experiments, which is one-time and negligible in distributed training initialization.
Based on Figure 11 and our experience, a practical guideline is to avoid excessively large FSDP group sizes and use hierarchical parallelism (e.g., HSDP) to scale to larger GPU counts. In practice, we select the FSDP group size by offline simulation to minimize LCM-induced rounding for a given model and RaggedShard granularity. When model dimensions remain configurable, choosing hidden sizes divisible by small composite factors rather than large co-prime dimensions can further reduce padding overhead.
@@ 6.5 Performance Breakdown
Table 2: Component ablation for 8-bit Adam, reported as normalized throughput after disabling each component independently. N/A indicates that the corresponding configuration is not meaningfully runnable without intrusive modifications to the model or optimizer code or manual management of custom collectives.
@@
  • DBuffer. Disabling DBuffer drops throughput by 7.2%, reflecting the Copy-In/Copy-Out overhead around collectives when communication buffers require copying.
@@
  • RaggedShard DTensor. Disabling RaggedShard makes it effectively non-runnable: users must either (i) carefully change every model and optimizer tensor so that $32\times 32$ block boundaries align with shard boundaries, or (ii) manually implement complex collectives (e.g., per-block metadata exchange and state gathering) to recover block-wise semantics. We therefore report this setting as N/A to indicate it is not meaningfully usable.
@@ 7 Lessons Learned
During the deployment of veScale-FSDP for real industrial workloads that use more than 10K GPUs, we summarize the key lessons we have learned.
@@
Lesson-2: Design system abstractions on the shoulders of giants. DTensor provides a powerful abstraction that already supports a wide range of parallelization techniques. By designing new abstractions on top of DTensor, we can seamlessly integrate existing parallelization strategies. In our work, RaggedShard is implemented as an optional placement on DTensor, enabling easy collaboration of established infrastructure such as Tensor and Expert Parallelism, as well as mature training tools like distributed checkpointing [dcp]. This approach minimizes engineering effort while contributing to a shared ecosystem that benefits the broader community. In fact, RaggedShard has already appeared as a planned feature on the official roadmap [pytorch_2026_h1_roadmap] of PyTorch.